(3.3.21) Use the following cell phone airport data speeds​ (Mbps) from a particular network. Find P20
0.1
0.1
0.2
0.2
0.3
0.5
0.5
0.5
0.6
0.6
0.6
0.7
0.7
0.9
0.9
1.1
1.2
1.2
1.6
1.9
2.2
2.3
2.4
2.7
2.7
3.3
3.5
3.6
3.7
4.5
5.4
5.7
5.8
6.4
7.6
10.1
10.8
11.3
11.5
11.8
11.8
12.5
13.2
13.5
13.8
14.1
14.3
15.1
15.5
29.9
P20=? Mbps
1
Expert's answer
2020-12-04T13:49:23-0500

k=20 and n=50 (n is amount of cell phone airport data speeds)(k is what P is)

L = (k/100) * n =(20/100) * 50 =10

now look up the 10th number and 11th number

P20 =(0.6 + 0.6)/2 = 0.6